Two companies, A and B, are worth $3 million and $6 million, respectively. A and B have CAPM expected rates of returns of 9% and 16.3%, respectively. If the two companies merge, what will the conglomerate's CAPM expected rates of return be? _____________________

Your production of widgets is governed by an average production cost equation, avg. cost = $1,000 + $5,000/(x+1000) where x is the number of goods produced. If you are currently selling 9,000 units at a price $3.44, what is the marginal cost of selling an additional unit?__________________________

Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

Company A = $3 million

Company B = $6 million

When the companies merge, there share in merged company is in the ratio 1:2 (as per their worth).

Company A share = 1/3

Company B share = 2/3

Conglomerate's CAPM expected rate of return = (1/3 x 9%) + (2/3 x 16.3%) = 3% + 10.87% = 13.87%